# Energy landscape analysis based on the Ising model: Tutorial review

### v1

@@ 2.3 Accuracy of fit
A least stringent test, which is often used, is to examine the correlation between $\langle\sigma_{i}\rangle_{\text{model}}$ and $\langle\sigma_{i}\rangle_{\text{data}}$ across the $i$ values and that between $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$ across $(i,j)$ pairs. A good fit of the PMEM to the data should yield $\langle\sigma_{i}\rangle_{\text{model}}\approx\langle\sigma_{i}\rangle_{\text{data}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}\approx\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$. However, these relationships are not enough for the following reason. If we run a gradient ascent algorithm to fit the PMEM (section 2.2), it will converge due to the concavity of the problem for any data while convergence may take a long time if $N$ is relatively large. Once the algorithm sufficiently converges, one obtains $\langle\sigma_{i}\rangle_{\text{model}}\approx\langle\sigma_{i}\rangle_{\text{data}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}\approx\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$ even if the data are a poor fit to the PMEM.
@@
is the Kullback-Leibler divergence between the probability distribution of the activity pattern in the $\ell$th-order $(\ell=1,2)$ maximum entropy model, $P_{\ell}(\bm{\sigma})$, and the distribution of the activity pattern in the original data, $P_{\text{data}}(\bm{\sigma})$. Distribution $P_{2}(\bm{\sigma})$ is equivalent to $P(\bm{\sigma})$ given by Eqs. (6) and (7). It is called second-order because the highest-order function in terms of $\sigma_{i}$ appearing in the model is of second order (i.e., $J_{ij}\sigma_{i}\sigma_{j}$). The first-order, or independent, maximum entropy model (i.e., $\ell=1$) is given by Eq. (6) but without interaction terms. In other words, we force $J_{ij}=0$ $\forall i,j$ in Eq. (7) and set each $h_{i}$ so that $\langle\sigma_{i}\rangle_{\text{model}}=\langle\sigma_{i}\rangle_{\text{data}}$ is satisfied. Quantity $r_{D}$ ranges between 00 and $1$ for the maximum likelihood estimator of the PMEM. We obtain $r_{D}=1$ when the PMEM perfectly fits the distribution of the activity pattern for the given data. In contrast, $r_{D}=0$ when the PMEM does not fit the data any better than the first-order model. For the maximum likelihood estimator, $r_{D}$ is equal to another often employed accuracy measure given by $(S_{1}-S_{2})/(S_{1}-S_{\text{data}})$, where $S_{\ell}=-\sum_{k=1}^{2^{N}}P_{\ell}(\bm{\sigma}^{(k)})\log P_{\ell}(\bm{\sigma}^{(k)})$ is the Shannon entropy of distribution $P_{\ell}$ [9, 11].
The accuracy of fit would be small when $N$ is large or $t_{\max}$ is small. In fact, the ELA is a data-hungry method. Therefore, it is important to check and report the value of $r_{D}$ or a different accuracy of fit measure. We recommend the value of $r_{D}$ above $0.85$, or ideally $0.9$, although these guideline values do not have any theoretical ground.
@@ 2.4 Local minima of energy and the network of activity patterns
Intuitively, a local minimum of energy is the bottom of a basin of the energy landscape shown in Figs. 1 and 2. However, the $2^{N}$ activity patterns do not lie in a one- or two-dimensional space as depicted in these figures. To precisely define the local minimum, let us consider Fig. 4, in which there are $N=5$ variables and an activity pattern $\bm{\sigma}^{(19)}=(-1,1,-1,-1,1)$ is neighbored by five other activity patterns. Each of these neighboring activity patterns is different from $\bm{\sigma}^{(19)}$ in the sign of just one $\sigma_{i}$. In other words, these five activity patterns are in Hamming distance $1$ from $\bm{\sigma}^{(19)}$. There are five activity patterns neighboring $\bm{\sigma}^{(19)}$ because one can flip any one entry of $\bm{\sigma}^{(19)}$ to obtain a neighboring activity pattern. Figure 4 shows that $\bm{\sigma}^{(19)}$ has a lower energy than all the five neighbors. Therefore, $\bm{\sigma}^{(19)}$ appears more frequently than each of its neighboring activity patterns in the data. When this is the case, we say that $\bm{\sigma}^{(19)}$ is a local minimum. If we allow the activity pattern to vary over time, the system’s state is more likely to transit from a neighbor of $\bm{\sigma}^{(19)}$, say $\bm{\sigma}^{(17)}$, to $\bm{\sigma}^{(19)}$ than vice versa because $\bm{\sigma}^{(19)}$ is more frequent than $\bm{\sigma}^{(17)}$. In this sense, $\bm{\sigma}^{(19)}$ is locally attractive.
@@
Figure 4: A schematic of the local minimum of energy. A system of $N=5$ variables is assumed. Activity pattern $\bm{\sigma}^{(19)}$ is a local minimum because it has a lower energy than all its neighbors. The neighbors are the activity patterns with Hamming distance 1 from $\bm{\sigma}^{(19)}$. The numbering scheme for the activity patterns is consistent with the one used in Fig. 3(c).
@@
In fact, any activity pattern has $N$ neighbors no matter whether or not it is a local minimum. The $2^{N}$ activity patterns form a “network of activity patterns” shown in Fig. 5. This network is called the hypercube. One can verify that each node of this network, corresponding to an activity pattern, has exactly $N$ neighbors. Using the terminology of network science (or graph theory), there are $2^{N}$ nodes, and the degree of each node (i.e., the number of edges the node has) is equal to $N$. Furthermore, in Fig. 5, each node has ten nodes in distance (i.e., minimum number of hops) 2, ten nodes in distance 3, five nodes in distance 4, and one node in distance 5. The single node in distance 5 is the one in which all the $N$ values of $\sigma_{i}$ are flipped. For example, the node in distance 5 from $\bm{\sigma}^{(19)}=(-1,1,-1,-1,1)$ is $\bm{\sigma}^{(14)}=(1,-1,1,1,-1)$. In general, there are $\binom{N}{d}$ nodes in distance $d$ from any node, where $\binom{}{}$ represents the binomial coefficient. The network of activity patterns represents admissible routes of the system’s state, $\bm{\sigma}(t)$, when $\bm{\sigma}(t)$ is allowed to change in one entry, $i$, at one time.
@@
Figure 5: Networks of activity patterns for $N\in\{1,\ldots,5\}$. They are hypercubes having $2^{N}$ nodes. Each node represents an activity pattern and has $N$ neighbors.
@@
In empirical data, $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$ are often different in more than one $i$ unless the time resolution of the observation is high. For example, in Fig. 3(b), $\bm{\sigma}(1)=(1,1,-1,-1,-1)$ and $\bm{\sigma}(2)=(1,-1,-1,1,-1)$ are different in two variables, i.e., $i=2$ and $i=4$. If $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$ are indeed different in more than one $i$ for some $t$, the trajectory is not a discrete-time walk on the hypercube because the walker can move from a node to its non-neighboring node in one time step. However, even in this situation, the assumption that only a single entry of $\bm{\sigma}(t)$ can flip at a time is still plausible by considering that we are observing a continuous-time walk in discrete time. This is the case if synchronous threshold crossing, which would simultaneously flip multiple $\sigma_{i}(t)$, is absent for the resolution of discrete times $t=1$, $2$, $\ldots$ at which one observes the continuous-time walk.
@@
To discuss the interpretation of $\bm{\sigma}(t)$ as a discrete-time observation of a continuous-time walk, we postulate an underlying $N$-dimensional dynamics in continuous time, $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$, and consider that we observe it in discrete times $t\in\{1,2,\ldots\}$, obtaining $\bm{\sigma}(t)$. Discrete time $t$ does not have to be the same as $t^{\prime}$. In Fig. 6, the $t$ shown in blue is identical with $t^{\prime}$. In contrast, the $t$ shown in red is sampled three times less frequently than the $t$ shown in blue. Recall that $\sigma_{i}(t)$ flips if and only if $x_{i}(t)$ crosses the threshold $\theta_{i}$, no matter whether the time is continuous or discrete. If different $x_{i}(t)$’s simultaneously cross their thresholds, then multiple $\sigma_{i}(t)$’s flip at the same time. In this case, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is not a nearest-neighbor walk on the hypercube. On the contrary, if such a synchronous threshold crossing is absent and the time resolution for sampling $\bm{\sigma}(t)$ is high enough, then $\bm{\sigma}(t)$ flips at most one bit at a time. In this case, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is a nearest-neighbor walk. In other words, the walker stays at one node of the hypercube for some discrete times, then moves to one of its neighbors, and so on, producing a discrete-time walk on the hypercube, without jumps from a node to a non-neighboring node. For example, the walk shown in blue in Fig. 6(a) is a nearest-neighbor walk. The figure indicates that $x_{j}(t^{\prime})$ changes from positive to negative first, followed by the same change in $x_{i}(t^{\prime})$. Then, $x_{k}(t^{\prime})$ changes from negative to positive. Then, $x_{i}(t^{\prime})$, $x_{k}(t^{\prime})$, and $x_{j}(t^{\prime})$ turn from positive to negative in this order. The discrete-time walk $\bm{\sigma}(t)$ shown in blue captures all these events including their chronology. In contrast, the walk shown in red in Fig. 6(a) is not a nearest-neighbor walk because the time resolution for observing $\bm{\sigma}(t)$ is too low. For example, both $\sigma_{1}(t)$ and $\sigma_{2}(t)$ change between $t=1$ and $t=2$.
@@
In the PMEM fitted to empirical data, we often find moves from a node on the hypercube to its non-nearest-neighbor, as shown in Fig. 3(b) and as $\bm{\sigma}(t)$ shown in red in Fig. 6(a). We justify this situation by regarding such $\bm{\sigma}(t)$ as an observation of originally continuously varying $\bm{x}(t^{\prime})$ at a relatively low time resolution, if there is no approximately simultaneous threshold crossing event.
@@
Figure 6(b) shows a case with approximately simultaneous threshold crossing events in the time series $\{\bm{x}(t^{\prime});t^{\prime}\in\mathbb{R}\}$, marked by the green and yellow circles. In this case, even if the time resolution for observing $\bm{\sigma}(t)$ is high, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is not a nearest-neighbor walk. For example, the walker shown in blue in Fig. 6(b) moves to a non-neighboring node between $t=2$ and $t=3$ and between $t=7$ and $t=8$, which are two intervals in which approximately simultaneous threshold crossing occurs. In fact, a move to a non-neighboring node also occurs between $t=6$ and $t=7$, which contains a less stringent simultaneous threshold crossing of $x_{j}(t^{\prime})$ and $x_{k}(t^{\prime})$. We also note that the same $\bm{\sigma}(t)$, shown in blue, fails to capture the two threshold crossings of $x_{k}(t^{\prime})$ between $t^{\prime}=t=8$ and $t^{\prime}=t=9$ and two threshold crossings of $x_{i}(t^{\prime})$ between $t^{\prime}=t=9$ and $t^{\prime}=t=10$. In this situation, it is difficult to justify the use of the PMEM and hence ELA.
@@
Approximately simultaneous threshold crossing is a relative concept. If we observe $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$, at large regular intervals to produce $\bm{\sigma}(t)$, then $x_{i}(t^{\prime})$ with multiple $i$’s would cross their thresholds, perhaps multiple times for some $i$’s, between adjacent discrete times (i.e., between $t$ and $t+1$). This is a case of simultaneous threshold crossing. In the extreme case, if the observation occurs with excessively long time intervals, then $\bm{x}(t)$ and $\bm{x}(t+1)$ would be completely uncorrelated, and so are $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$. Then, the trajectory $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ will be composed of independent samples drawn from the distribution $P(\bm{\sigma})$, completely ignoring the hypercube structure. Given this discussion, we could try to justify the use of the PMEM and ELA even in the case of $\bm{\sigma}(t)$ shown in blue and red in Fig. 6(b). If we use a higher sampling rate, such as $t=100t^{\prime}$, such that $t\in\{1,2,\ldots\}$ corresponds to $t^{\prime}\in\{0.01,0.02,\ldots\}$, then $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ will be a nearest-neighbor walk unless there are perfectly simultaneous threshold crossing events in $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$. However, excessively synchronous threshold crossing of $x_{i}(t^{\prime})$’s would violate the detailed balance condition underlying the random walk assumption (section 3.2.3). Therefore, we do not recommend that users rely on random-walk views when strongly synchronous dynamics among the variables is known or expected.
@@
Figure 6: Time series of activity pattern as a discrete-time observation of a continuous-time walk. (a) Absence and (b) presence of synchronous threshold crossing. The horizontal dotted lines represent the threshold for binarization, $\theta_{i}$. We show the case $t=t^{\prime}$ and $t=3t^{\prime}-2$ in blue and red, respectively. In (b), the green circles show an approximately simultaneous threshold crossing of $x_{i}(t^{\prime})$, $x_{j}(t^{\prime})$, and $x_{k}(t^{\prime})$. The yellow circles show an approximately simultaneous threshold crossing of $x_{i}(t^{\prime})$ and $x_{j}(t^{\prime})$.

### v2

@@ 2.3 Accuracy of fit
A least stringent test, which is often used, is to examine the correlation between $\langle\sigma_{i}\rangle_{\text{model}}$ and $\langle\sigma_{i}\rangle_{\text{data}}$ across the $i$ values and that between $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$ across $(i,j)$ pairs. A good fit of the PMEM to the data should yield $\langle\sigma_{i}\rangle_{\text{model}}\approx\langle\sigma_{i}\rangle_{\text{data}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}\approx\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$, where $\approx$ represents “approximately equal to”. However, these relationships are not enough for the following reason. If we run a gradient ascent algorithm to fit the PMEM (Section 2.2), it will converge due to the concavity of the problem for any data while convergence may take a long time if $N$ is relatively large. Once the algorithm sufficiently converges, one obtains $\langle\sigma_{i}\rangle_{\text{model}}\approx\langle\sigma_{i}\rangle_{\text{data}}$ and $\langle\sigma_{i}\sigma_{j}\rangle_{\text{model}}\approx\langle\sigma_{i}\sigma_{j}\rangle_{\text{data}}$ even if the data are a poor fit to the PMEM.
@@
is the Kullback-Leibler divergence between the probability distribution of the activity pattern in the $\ell$th-order $(\ell=1,2)$ maximum entropy model, $P_{\ell}(\bm{\sigma})$, and the distribution of the activity pattern in the original data, $P_{\text{data}}(\bm{\sigma})$. Distribution $P_{2}(\bm{\sigma})$ is equivalent to $P(\bm{\sigma})$ given by Eqs (6) and (7). It is called second-order because the highest-order function in terms of $\sigma_{i}$ appearing in the model is of second order (i.e., $J_{ij}\sigma_{i}\sigma_{j}$). The first-order, or independent, maximum entropy model (i.e., $\ell=1$) is given by Eq (6) but without interaction terms. In other words, we force $J_{ij}=0$ $\forall i,j$ in Eq (7) and set each $h_{i}$ so that $\langle\sigma_{i}\rangle_{\text{model}}=\langle\sigma_{i}\rangle_{\text{data}}$ is satisfied. Quantity $r_{D}$ ranges between 00 and $1$ for the maximum likelihood estimator of the PMEM. We obtain $r_{D}=1$ when the PMEM perfectly fits the distribution of the activity pattern for the given data. In contrast, $r_{D}=0$ when the PMEM does not fit the data any better than the first-order model. For the maximum likelihood estimator, $r_{D}$ is equal to another often employed accuracy measure given by $(S_{1}-S_{2})/(S_{1}-S_{\text{data}})$, where $S_{\ell}=-\sum_{k=1}^{2^{N}}P_{\ell}(\bm{\sigma}^{(k)})\log P_{\ell}(\bm{\sigma}^{(k)})$ is the Shannon entropy of distribution $P_{\ell}$ [9, 11].
The accuracy of fit would be small when $N$ is large or $t_{\max}$ is small. In fact, the ELA is a data-hungry method. Therefore, it is important to check and report the value of $r_{D}$ or a different accuracy of fit measure. We recommend the value of $r_{D}$ above $0.85$, or ideally $0.9$. This is because, although these guideline values do not have any theoretical ground, an influential early study of the PMEM applied to neural spiking data realized $r_{D}\approx 0.9$ [7].
@@ 2.4 Local minima of energy and the network of activity patterns
Intuitively, a local minimum of energy is the bottom of a basin of the energy landscape shown in Figs 1 and 2. However, the $2^{N}$ activity patterns do not lie in a one- or two-dimensional space as depicted in these figures. To precisely define the local minimum, let us consider Fig 4, in which there are $N=5$ variables and an activity pattern $\bm{\sigma}^{(19)}=(-1,1,-1,-1,1)$ is neighbored by five other activity patterns. Each of these neighboring activity patterns is different from $\bm{\sigma}^{(19)}$ in the sign of just one $\sigma_{i}$. In other words, these five activity patterns are in Hamming distance $1$ from $\bm{\sigma}^{(19)}$. There are five activity patterns neighboring $\bm{\sigma}^{(19)}$ because one can flip any one entry of $\bm{\sigma}^{(19)}$ to obtain a neighboring activity pattern. Fig 4 shows that $\bm{\sigma}^{(19)}$ has a lower energy than all the five neighbors. Therefore, $\bm{\sigma}^{(19)}$ appears more frequently than each of its neighboring activity patterns in the data. When this is the case, we say that $\bm{\sigma}^{(19)}$ is a local minimum. If we allow the activity pattern to vary over time, the system’s state is more likely to transit from a neighbor of $\bm{\sigma}^{(19)}$, say $\bm{\sigma}^{(17)}$, to $\bm{\sigma}^{(19)}$ than vice versa because $\bm{\sigma}^{(19)}$ is more frequent than $\bm{\sigma}^{(17)}$. In this sense, $\bm{\sigma}^{(19)}$ is locally attractive.
@@
Fig 4: A schematic of the local minimum of energy. A system of $N=5$ variables is assumed. Activity pattern $\bm{\sigma}^{(19)}$ is a local minimum because it has a lower energy than all its neighbors. The neighbors are the activity patterns with Hamming distance 1 from $\bm{\sigma}^{(19)}$. The numbering scheme for the activity patterns is consistent with the one used in Fig 3(c).
@@
In fact, any activity pattern has $N$ neighbors no matter whether or not it is a local minimum. The $2^{N}$ activity patterns form a “network of activity patterns” shown in Fig 5. This network is called the hypercube. One can verify that each node of this network, corresponding to an activity pattern, has exactly $N$ neighbors. Using the terminology of network science (or graph theory), there are $2^{N}$ nodes, and the degree of each node (i.e., the number of edges the node has) is equal to $N$. Furthermore, in Fig 5, each node has ten nodes in distance (i.e., minimum number of hops) 2, ten nodes in distance 3, five nodes in distance 4, and one node in distance 5. The single node in distance 5 is the one in which all the $N$ values of $\sigma_{i}$ are flipped. For example, the node in distance 5 from $\bm{\sigma}^{(19)}=(-1,1,-1,-1,1)$ is $\bm{\sigma}^{(14)}=(1,-1,1,1,-1)$. In general, there are $\binom{N}{d}$ nodes in distance $d$ from any node, where $\binom{}{}$ represents the binomial coefficient. The network of activity patterns represents admissible routes of the system’s state, $\bm{\sigma}(t)$, when $\bm{\sigma}(t)$ is allowed to change in one entry, $i$, at one time.
@@
Fig 5: Networks of activity patterns for $N\in\{1,\ldots,5\}$. They are hypercubes having $2^{N}$ nodes. Each node represents an activity pattern and has $N$ neighbors.
@@
In empirical data, $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$ are often different in more than one $i$ unless the time resolution of the observation is high. For example, in Fig 3(b), $\bm{\sigma}(1)=(1,1,-1,-1,-1)$ and $\bm{\sigma}(2)=(1,-1,-1,1,-1)$ are different in two variables, i.e., $i=2$ and $i=4$. If $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$ are indeed different in more than one $i$ for some $t$, the trajectory is not a discrete-time walk on the hypercube because the walker can move from a node to its non-neighboring node in one time step. However, even in this situation, the assumption that only a single entry of $\bm{\sigma}(t)$ can flip at a time is still plausible by considering that we are observing a continuous-time walk in discrete time. This is the case if synchronous threshold crossing, which would simultaneously flip multiple $\sigma_{i}(t)$, is absent for the resolution of discrete times $t=1$, $2$, $\ldots$ at which one observes the continuous-time walk.
@@
To discuss the interpretation of $\bm{\sigma}(t)$ as a discrete-time observation of a continuous-time walk, we postulate an underlying $N$-dimensional dynamics in continuous time, $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$, and consider that we observe it in discrete times $t\in\{1,2,\ldots\}$, obtaining $\bm{\sigma}(t)$. Discrete time $t$ does not have to be the same as $t^{\prime}$. In Fig 6, the $t$ shown in blue is identical with $t^{\prime}$. In contrast, the $t$ shown in red is sampled three times less frequently than the $t$ shown in blue. Recall that $\sigma_{i}(t)$ flips if and only if $x_{i}(t)$ crosses the threshold $\theta_{i}$, no matter whether the time is continuous or discrete. If different $x_{i}(t)$’s simultaneously cross their thresholds, then multiple $\sigma_{i}(t)$’s flip at the same time. In this case, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is not a nearest-neighbor walk on the hypercube. On the contrary, if such a synchronous threshold crossing is absent and the time resolution for sampling $\bm{\sigma}(t)$ is high enough, then $\bm{\sigma}(t)$ flips at most one bit at a time. In this case, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is a nearest-neighbor walk. In other words, the walker stays at one node of the hypercube for some discrete times, then moves to one of its neighbors, and so on, producing a discrete-time walk on the hypercube, without jumps from a node to a non-neighboring node. For example, the walk shown in blue in Fig 6(a) is a nearest-neighbor walk. The figure indicates that $x_{j}(t^{\prime})$ changes from positive to negative first, followed by the same change in $x_{i}(t^{\prime})$. Then, $x_{k}(t^{\prime})$ changes from negative to positive. Then, $x_{i}(t^{\prime})$, $x_{k}(t^{\prime})$, and $x_{j}(t^{\prime})$ turn from positive to negative in this order. The discrete-time walk $\bm{\sigma}(t)$ shown in blue captures all these events including their chronology. In contrast, the walk shown in red in Fig 6(a) is not a nearest-neighbor walk because the time resolution for observing $\bm{\sigma}(t)$ is too low. For example, both $\sigma_{1}(t)$ and $\sigma_{2}(t)$ change between $t=1$ and $t=2$.
@@
In the PMEM fitted to empirical data, we often find moves from a node on the hypercube to its non-nearest-neighbor, as shown in Fig 3(b) and as $\bm{\sigma}(t)$ shown in red in Fig 6(a). We justify this situation by regarding such $\bm{\sigma}(t)$ as an observation of originally continuously varying $\bm{x}(t^{\prime})$ at a relatively low time resolution, if there is no approximately simultaneous threshold crossing event.
@@
Fig 6(b) shows a case with approximately simultaneous threshold crossing events in the time series $\{\bm{x}(t^{\prime});t^{\prime}\in\mathbb{R}\}$, marked by the green and yellow circles. In this case, even if the time resolution for observing $\bm{\sigma}(t)$ is high, $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ is not a nearest-neighbor walk. For example, the walker shown in blue in Fig 6(b) moves to a non-neighboring node between $t=2$ and $t=3$ and between $t=7$ and $t=8$, which are two intervals in which approximately simultaneous threshold crossing occurs. In fact, a move to a non-neighboring node also occurs between $t=6$ and $t=7$, which contains a less stringent simultaneous threshold crossing of $x_{j}(t^{\prime})$ and $x_{k}(t^{\prime})$. We also note that the same $\bm{\sigma}(t)$, shown in blue, fails to capture the two threshold crossings of $x_{k}(t^{\prime})$ between $t^{\prime}=t=8$ and $t^{\prime}=t=9$ and two threshold crossings of $x_{i}(t^{\prime})$ between $t^{\prime}=t=9$ and $t^{\prime}=t=10$. In this situation, it is difficult to justify the use of the PMEM and hence ELA.
@@
Approximately simultaneous threshold crossing is a relative concept. If we observe $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$, at large regular intervals to produce $\bm{\sigma}(t)$, then $x_{i}(t^{\prime})$ with multiple $i$’s would cross their thresholds, perhaps multiple times for some $i$’s, between adjacent discrete times (i.e., between $t$ and $t+1$). This is a case of simultaneous threshold crossing. In the extreme case, if the observation occurs with excessively long time intervals, then $\bm{x}(t)$ and $\bm{x}(t+1)$ would be completely uncorrelated, and so are $\bm{\sigma}(t)$ and $\bm{\sigma}(t+1)$. Then, the trajectory $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ will be composed of independent samples drawn from the distribution $P(\bm{\sigma})$, completely ignoring the hypercube structure. Given this discussion, we could try to justify the use of the PMEM and ELA even in the case of $\bm{\sigma}(t)$ shown in blue and red in Fig 6(b). If we use a higher sampling rate, such as $t=100t^{\prime}$, such that $t\in\{1,2,\ldots\}$ corresponds to $t^{\prime}\in\{0.01,0.02,\ldots\}$, then $\{\bm{\sigma}(1),\bm{\sigma}(2),\ldots\}$ will be a nearest-neighbor walk unless there are perfectly simultaneous threshold crossing events in $\bm{x}(t^{\prime})$, $t^{\prime}\in\mathbb{R}$. However, excessively synchronous threshold crossing of $x_{i}(t^{\prime})$’s would violate the detailed balance condition underlying the random walk assumption (Section 3.2.3). Therefore, we do not recommend that users rely on random-walk views when strongly synchronous dynamics among the variables is known or expected.
@@
Fig 6: Time series of activity pattern as a discrete-time observation of a continuous-time walk. (a) Absence and (b) presence of synchronous threshold crossing. The horizontal dotted lines represent the threshold for binarization, $\theta_{i}$. We show the case $t=t^{\prime}$ and $t=3t^{\prime}-2$ in blue and red, respectively. In (b), the green circles show an approximately simultaneous threshold crossing of $x_{i}(t^{\prime})$, $x_{j}(t^{\prime})$, and $x_{k}(t^{\prime})$. The yellow circles show an approximately simultaneous threshold crossing of $x_{i}(t^{\prime})$ and $x_{j}(t^{\prime})$.